# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v8

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph.
In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers.
This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes.
This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node.
Further, our approach finds the occurrence of dense regular graphs inside containers.
This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property.
Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
@@ 1 Introduction
Despite decades of research, the SSNC remains one of graph theory’s challenging open problems.
Special cases of the conjecture have been explored extensively.
Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six.
Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
Chen et al. [3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture.
This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
The concept of $m$-free graphs is particularly relevant to this paper, as it closely aligns with the work of Daamouch [7]. A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$. In his paper, Daamouch introduces induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9].
Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers.
This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC.
Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design.
For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node.
We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems.
The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
@@ 2 Graph Theory Terminology
$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
@@
Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
@@
We will focus our attention on Conjecture 1.1 and Conjecture 2.1. Conjecture 1.1 is easier to work with initially and understand why a node’s degree will double in the square graph. Concentrating on neighborhood analysis yielded numerous insights and features that aided in the development of our algorithm.
@@ 3 Approach and Motivation
We begin our analysis by considering generally oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes or arcs are being removed to form these partitions. This is a conceptual idea. Instead, think of it as putting a box or circle around the nodes and arcs. Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another and will have to reconnect to solve the global problem. However, thinking of them in this way allows us to use their independent nature.
This represents a stark contrast to most previous approaches to the SSNC. Many of these previous approaches were conceptual or theoretical. Papers would mention the existence of a node whose degree doubled. The notion, though, of a path to that node was not part of the paper. Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node. That presents a problem because oriented graphs can be varying in nature. This is the first global search technique on the SSNC problem.
@@ 4 Initial Lemmas
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making it at least as large as $v_{0}$’s first neighborhood.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence $v_{0}$ will have at least two second neighbors—$v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first. Thus, its out-degree will double in $G^{2}$.
@@
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, to construct our counterexample, we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
@@
We know that $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$. This means that $v_{0}$ has at least three second neighbors. As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$. ∎
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
@@
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{1}$ will double. See lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1)
@@
We did not require the ”right” starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimal degree node due to its ability to split the graph’s nodes.
@@ 5 Building a Data Structure
In this section, we develop a new data structure for oriented graphs (Graph Level Order). This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond. In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as containers. Additionally, we shall discuss interior, exterior, and back arcs and show the influence they have on our data structure. Finally, we will speak about how this data structure can use the divide and conquer paradigm to help solve the SSNC.
@@
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from $v_{0}$, with no assumptions on container relationships or sizes.
@@
$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where
$V(G)$ denotes the graph $G$’s vertex set, and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
@@
We will investigate the structure of oriented graphs from two angles—from the viewpoint of containers or $v_{0}$. Containers are analyzed based on their distance from $v_{0}$. Nodes are seen based on their position within these containers. They are either before a container, in a container, or after a container.
@@
In particular, we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the component $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to $v_{0}$. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes. We will speak more about back arcs in the next section.
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Next we will prove that any oriented graph can be represented in a GLOVER. We will need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties. This includes maintaining a total order of the containers, allowing for the comparison of any two nodes, establishing a total order of those nodes within the containers, and establishing interior and exterior neighbors.
@@
Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
  • Total Order on Elements within $G$ Nodes are partitioned based on their distance from $v_{0}$. This is a total order on the nodes of $G$. For any two nodes $u$ and $v$, this total order is determined by their containers.
@@
  • Comparability within Containers—For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
  • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
@@
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors.
    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
@@
What the GLOVER data structure does is transform the containers and distance metric into an ordering for containers. This ordering is more than just representative, though. The containers help to organize the nodes. What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from $v_{0}$. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering. However, it is a bold assumption to claim that oriented graphs have no back arcs. Oriented graphs come in various types; surely some exist with back arcs.
This is an important claim to investigate, and the next section will look into it in depth. We will prove that these back arcs do not hurt the algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node.
Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering. Mathematical induction is a proof technique that uses inference to prove statements. Remember that a well ordering is a concept in discrete mathematics where a total ordering on a set $S$ has the property that every non-empty subset of $S$ has a least element.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
@@
A discrete total order is a total order where each element has a distinct ”next” element in the order.
Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered.
@@
We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from $v_{0}$.

### v9

@@ Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes. This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node. Further, our approach finds the occurrence of dense regular graphs inside containers. This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property. Beyond theoretical implications, the algorithm and data structure have practical applications in data science, network optimization and algorithm design.
Keywords: degree doubling, oriented graph, square, algorithm
@@ 1 Introduction
Despite decades of research, the SSNC has remained one of graph theory’s challenging open problems. Special cases of the conjecture have been explored extensively. Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six. Brantner et al. [2] demonstrated that if the conjecture were false, infinitely many counterexamples would exist. Chen et al. [4] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $\lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [6] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [3], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
The concept of $m$-free graphs is particularly relevant to this paper, as our work closely aligns with the work of Daamouch [1]. A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$. In his paper, Daamouch works with the concept of induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9]. Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC. Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, algorithm design, and the development of data-driven methods for networked systems, including A/B testing. For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections matches or surpasses their direct contacts, offering insights into the spread of infections. In the development of algorithms for A/B testing within networked environments, understanding influence propagation is critical for designing methods that account for dependencies between individuals, ultimately leading to more accurate assessments of interventions.
Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
@@ 2 Graph Theory Terminology
$E^{2}=\{(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E\}$
@@
Definition 2.2.
Let $G=(V,E)$, the first neighborhood of a vertex $v\in V$ is denoted
$N^{+}(v)=\{w\in V(G)|(v,w)\in E)\}$
Definition 2.3.
Let $G=(V,E)$, the second neighborhood of a vertex $v\in V$ is denoted
$N^{++}(v)=\{u\in V(G)|(v,w)\text{ and }(w,u)\in E)\}$
@@
Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when the oriented graph is squared.
@@
We will focus our attention on both conjectures 1.1 and 2.1. Conjecture 1.1 is easier to work with initially and understand why a node’s degree will double in the square graph. Concentrating on neighborhood analysis yielded numerous insights and features that aided in the development of our algorithm. This is because it is not necessary to work with a different graph to understand degree-doubling properties. However, conjecture 2.1 could have just as easily been used and in some of the later proofs we will see this.
@@ 3 Approach and Motivation
We begin our analysis by considering generally oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $\delta$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we can now partition the graph into $\delta+1$ subgraphs corresponding to $\delta+1=d^{+}(v_{0})+1$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes or arcs are being removed from the graph to form these partitions. This is a conceptual idea. Instead, think of it as putting a box or circle around the nodes and arcs. Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another and will have to reconnect to solve the global problem. However, thinking of them in this way allows us to use their independent nature.
This represents a stark contrast to most previous approaches to the SSNC. Many of these previous approaches were conceptual or theoretical. Papers would mention the existence of a node whose degree doubled. The notion, though, of a path to that node was not part of the paper. Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node. That presents a problem because oriented graphs can be varying in nature. This may be the first global search technique on the SSNC problem.
@@ 4 Initial Lemmas
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making $v_{0}$’s second neighborhood at least as large as its first neighborhood.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. Because this is an oriented graph, the node $v_{0}$ cannot be one of these nodes. Hence $v_{0}$ will have at least two second neighbors—$v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first. Thus, its out-degree will double in $G^{2}$.
@@
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, to construct our counterexample, we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum degree node has very few out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
@@
This means that $v_{0}$ has at least three second neighbors. As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$. ∎
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors. By assumption, $v_{1}$ must have at least as many first neighbors as $v_{0}$.
@@
Figure 3: Illustration of Lemma 3: Minimum degree node v0 (degree 3) and its neighbors (degree 1 in G[N+(v0)], degree 2 elsewhere). In G2, v1’s degree doubles. See lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1)
@@
We did not require the right starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimal degree node due to its ability to split the graph’s nodes.
@@ 5 Building a Data Structure
5.1 Definitions
In this section, we develop a new data structure for oriented graphs Graph Level Order (GLOVER). This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond. In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as containers. Additionally, we shall discuss interior, exterior, and back arcs and show the influence they have on our data structure. Finally, we will speak about how this data structure can use the divide and conquer paradigm to help solve the SSNC.
@@ 5.1 Definitions
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node $v_{0}$, with no assumptions on container relationships or sizes.
@@
$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and dist}(v_{0},u)=\text{dist}(v_{0},v)=i\}$
where $V(G)$ denotes the graph $G$’s vertex set, and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
@@
We will investigate the structure of oriented graphs from two angles—from the viewpoint of containers and the minimum degree node $v_{0}$. Containers will be analyzed based on their distance from $v_{0}$. Nodes are seen based on their position within these containers. They are either before a container, in a container, or after a container.
@@
In particular, we saw things about interior neighbors in lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the container $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to $v_{0}$. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes. We will speak more about back arcs in section 8 (Back Arcs).
5.2 Graph Level Order
@@ A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Total Order on Elements of G: We need that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
@@
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors if $(x,y)\in G$.
  • If $x\in C_{i}$ and $y\in C_{i+1}$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Next, we will prove that any oriented graph can be represented in a GLOVER. We will need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties. This includes maintaining a total order of the containers, allowing for the comparison of any two nodes, establishing a total order of those nodes within the containers, and establishing interior and exterior neighbors.
@@
Each of these containers $C_{i}=\{u|\text{dist}(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$.
@@
  • Total Order on Elements within $G$—Nodes are partitioned based on their distance from $v_{0}$. We know that distance is a total order on the nodes of $G$.
    To prove the containers are a total order, we need to show that any two nodes in $G$ can be compared by their container. Let $u,v$ be two nodes in $G$ such that $u\neq v$. We assume that both $u$ and $v$ are reachable from $v_{0}$.
    First, we consider the situation where $u$ is in a closer container than $v$. This implies that $\text{dist}(v_{0},u)<\text{dist}(v_{0},v)$.
    Next, we will look at the situation where $u$ and $v$ are in the same container. This implies that $\text{dist}(v_{0},u)=\text{dist}(v_{0},v)$.
    Finally, we consider the situation where $u$ is in a further container from $v_{0}$ than $v$. This implies that $\text{dist}(v_{0},u)>\text{dist}(v_{0},v)$
    In each situation we were able to translate the container relationship to a distance relationship. We know that distance is already a total order. This shows that the container relationship is also a total order.
  • Comparability within Containers—For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric (degree) to distinguish $u$ from $v$ or compare $u$ to $v$.
  • Consistency of Comparisons across Containers—Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
@@
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors if there is an arc $(u,v)\in G$
    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1 and if there is an arc $(u,v)\in G$
@@
What the GLOVER data structure does is transform the oriented graph into an ordered of containers with metrics. This ordering is more than just representative, though. The containers help to organize the nodes. What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from $v_{0}$. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$. Now, each node has a proper place inside the container, up to reordering.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering. However, it is a bold assumption to claim that oriented graphs have no back arcs. Oriented graphs come in all shapes and sizes; surely some exist with back arcs.
This is an important claim to investigate, and section 8 (Back Arcs) will look into it in depth. We will prove that when these back arcs do exist, we are not able to construct a GLOVER data structure. Instead, we are able to immediately find a degree-doubling node. That is, a back arc immediately leads us to a node whose second neighborhood is at least as large as its first neighborhood. This means that back arcs do not hurt with the development of our algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node(s).
Next, we will show that one of the other benefits of this data structure is mathematical induction via well ordering. Mathematical induction is a proof technique that uses inference to prove statements. Remember that a well ordering is a concept in discrete mathematics where a total ordering on a set $S$ has the property that every non-empty subset of $S$ has a least element.
@@
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order.
@@
A discrete total order is a total order where each element has a distinct next element in the order.
Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a smallest element according to the order. This fulfills the condition of being well-ordered.
@@
We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from $v_{0}$-the minimum degree node.